OpenStudy (livya15):

**Will give MEDAL** A telephone company finds there is a net profit of $15 per instrument if a given network has 1000 users. For every user over 1000, the profit decreases by 1 cent. What is the maximum profit the company can make per network?

OpenStudy (livya15):

@Hero

jimthompson5910 (jim_thompson5910):

n = number of users p = profit per user Currently n = 1000 p = 15 so the total profit would be T = n*p T = 1000*15 T = 15000 Let x = number of additional users added If we add on x users, then we go from n = 1000 to n = 1000+x For each user added, we lose 1 cent ($0.01) in profit, per user, so p = 15 - 0.01x The new total profit T is T = n*p T = (1000+x)*(15-0.01x) Your goal is to find the vertex of the new T function. That will allow you to determine the maximum profit possible.

jimthompson5910 (jim_thompson5910):

Hint: I recommend expanding out `(1000+x)*(15-0.01x)` as your first step

OpenStudy (livya15):

I expanded that equation and got: T=15000+5x-.01x^2

OpenStudy (livya15):

Then i took the derivative and got: T'=5-.02x When I set 0=5-.02x, I got x=250

jimthompson5910 (jim_thompson5910):

very good

jimthompson5910 (jim_thompson5910):

x = 250 will occur at the max. Plug in x = 250 to find the value of T

OpenStudy (livya15):

I plugged the x=250 into the original equation and got the answer of $15625. Does this sound like the correct answer?

jimthompson5910 (jim_thompson5910):

yep, the max profit possible is $15625 so if you add on 250 more users, then you gain an additional $625 (15625-15000=625)

OpenStudy (livya15):

Thank you!!!

jimthompson5910 (jim_thompson5910):

you're welcome